# Probabilistic Analysis of Copyright Disputes and Generative AI Safety

### v4

@@ Abstract.
This paper presents a probabilistic approach to analyzing copyright infringement disputes. Under this approach, evidentiary principles shaped by case law are formalized in probabilistic terms, allowing for a mathematical examination of issues arising in such disputes. The usefulness of this approach is showcased through its application to the “inverse ratio rule”—a controversial legal doctrine adopted by some courts. Although this rule has faced significant criticism, a formal proof demonstrates its validity, provided it is properly defined. Furthermore, the paper employs the probabilistic approach to study the copyright safety of generative AI. Specifically, the Near Access-Free (NAF) condition, previously proposed as a strategy for mitigating the heightened copyright infringement risks of generative AI, is evaluated. The analysis reveals that, while the NAF condition mitigates some infringement risks, its justifiability and efficacy are questionable in certain contexts. These findings illustrate how taking a probabilistic perspective can enhance our understanding of copyright jurisprudence and its interaction with generative AI technology.
@@ 1. Introduction
This article employs a probabilistic approach to analyze copyright infringement disputes, seeking to bring clarity to its complexities. The analysis is carried out by distilling and formalizing the fundamental principles governing copyright disputes as established through case law in the United States, particularly focusing on the requirements a copyright owner must meet to make a prima facie case of copyright infringement in court. This body of jurisprudence encompasses not only the substantive elements of copyright law, but also the evidentiary principles that guide how evidence is presented and assessed (Loren and Reese, 2019).
@@
The theoretical foundation of this work lies in the long-recognized link between probabilistic reasoning and judicial proof (Eggleston, 1989; Cullison, 1969; Broun and Kelly, 1970; Finkelstein and Fairley, 1970; Cohen, 1977; Eggleston, 1978; McCauliff, 1982). Specifically, the legal fact-finding process can be understood to involve assessment of the probabilities of allegations being true or false based on the available evidence (Stein, 2005a; Cheng, 2013). Building on this link, core evidentiary principles in copyright jurisprudence can be formalized in probabilistic terms, allowing for a mathematical analysis of a controversial legal doctrine, the “inverse ratio rule.” This rule governs the relationship between two critical pieces of evidence in copyright infringement cases—access and similarity to the original work. Although the validity of this rule has been the subject of considerable debate (Aronoff, 2007; Balganesh and Menell, 2022), a formal proof is established here, demonstrating that it holds under natural assumptions when its content and scope are carefully delineated.
Furthermore, this approach facilitates the analysis of copyright infringement risks introduced by technological advancements and the identification of effective mitigation strategies. Specifically, the probabilistic approach is utilized to evaluate the efficacy of the Near Access-Free (NAF) condition (Vyas et al., 2023), which has been proposed to mitigate the risk of copyright infringement posed by generative AI. This paper identifies extensive access to copyrighted materials by generative models as a significant source of heightened infringement risks. While the NAF condition addresses this issue to some extent, the analysis reveals that it does so partly on ethically questionable grounds—namely, by reducing transparency regarding the materials to which generative AI had access during training—and that its effectiveness in mitigating infringement risks remains limited under certain situations.
2. Related Work
Existing research has investigated the relationship between copyright law and probabilistic thinking, specifically concerning how the probability of creation relates to concepts of creativity and originality within the context of copyright law (Byron, 2006; Khong, 2007; Vermont, 2012; Hacohen and Elkin-Koren, 2024; Chiba-Okabe and Su, 2024). However, there has been little discussion on the application of probabilistic approaches in analyzing the evidentiary rules of copyright infringement disputes. While a previous study has alluded to the connection between the probabilistic nature of legal fact-finding and the inverse ratio rule (Balganesh and Menell, 2022), no formal probabilistic analysis of the rule has been undertaken, to the best of the author’s knowledge. This gap is addressed by employing a probabilistic approach to formalize the evidentiary principles governing copyright disputes and mathematically analyze the inverse ratio rule.
In response to rapid advancements in generative AI technology, a substantial body of literature has emerged, addressing potential copyright risks and defenses available to developers (Sag, 2023; Henderson et al., 2023; Lee et al., 2024). These works discuss various legal challenges, including the potential for copyright infringement arising from AI-generated content, with some studies analyzing the Near Access-Free (NAF) condition as a mitigation strategy (Lee et al., 2024; Elkin-Koren et al., 2024). This paper contributes to the discourse by providing a probabilistic analysis of legal issues surrounding the NAF condition, particularly through the lens of inference based on evidence in copyright infringement disputes.
3. Analysis of Copyright Disputes
3.1. Essential Evidentiary Principles in Copyright Disputes
To prevail in a copyright infringement dispute, the copyright owner, as the plaintiff, must demonstrate two elements under the traditional Arnstein test: “(a) that defendant copied from plaintiff’s copyrighted work and (b) that the copying (assuming it to be proved) went so far as to constitute improper appropriation.” [^1] Arnstein v. Porter, 154 F.2d 464, 468 (2d Cir. 1946). These two requirements are often referred to simply as “factual copying” and “improper appropriation,” respectively (Balganesh and Menell, 2022). Although several defenses, including fair use, may be available to the defendant in some cases even if these elements are established (Loren and Reese, 2019), these defenses are not considered in this paper to maintain focus on the essential structure of copyright infringement cases.
Direct evidence of factual copying includes admission by the defendant or an eyewitness account of the replication or appropriative behavior, but such evidence is hard to obtain. [^2] See, e.g., Smith v. Jackson, 84 F.3d 1213, 1218 (9th Cir. 1996), stating “direct evidence of copying is not available in most cases.” Therefore, factual copying is almost invariably proved by circumstantial evidence (Gorman, 2006; Balganesh and Menell, 2022), and the evidence generally accepted by courts consists of that pertaining to access and similarity. Specifically, copying is proved based on (i) evidence showing that the defendant had access to the original work and (ii) the similarity between the original work and the defendant’s work (Balganesh and Menell, 2022; Loren and Reese, 2019). Courts generally hold that proof of access combined with a certain level of similarity, that is probative of copying, is sufficient to establish factual copying (Latman, 1990; Loren and Reese, 2019). Although both access and similarity support the inference of factual copying, simply having proof of access does not suffice to establish factual copying when there is no similarity between the two works. [^3] Arnstein, 154 F.2d at 468. Proof of access pertains to the antecedent of factual copying, in the sense that the defendant having had access to the original work is a necessary condition for engaging in copying (Balganesh and Menell, 2022; Latman, 1990). In contrast, similarity emerges as a result of copying (Balganesh and Menell, 2022; Latman, 1990), and when the similarity between the two works is so pronounced as to be deemed “striking,” this alone suffices to establish that copying has occurred. [^4] Id. at 468-69. In such cases, it can be inferred that the defendant had access to the original work based on the high degree of similarity. [^5] Ty, Inc. v. GMA Accessories, Inc., 132 F.3d 1167, 1170 (7th Cir. 1997), stating “a similarity that is so close as to be highly unlikely to have been an accident of independent creation is evidence of access.”
The second requirement, improper appropriation, is established by demonstrating that the extent of copying reaches the threshold at which the similarity between the original work and the defendant’s work can be deemed “substantial” in degree. Determination of this requirement involves a normative judgment, taking into account both quantitative and qualitative aspects of copying (Gorman, 2006; Balganesh, 2012). Thus, similarity serves a dual role: it supports the finding of factual copying and acts as an independent criterion for establishing improper appropriation (Latman, 1990; Balganesh and Menell, 2022), though the sense of similarity differs in each context, even if there may be some overlap (Loren and Reese, 2019).
@@ 3.2. Formalization of Evidentiary Principles
Legal fact-finding is a reasoning process that involves assessing and weighing available evidence to reach a decision amid uncertainty. This process may be appropriately characterized by probabilistic reasoning, wherein adjudicators employ common sense and logic, drawing on knowledge of the frequencies of occurrences of certain events and patterns derived from experience to evaluate the likelihood of the particular circumstances in question (Stein, 2005b). Although the absolute truth typically cannot be attained, the final decision in most civil cases is made based on the standard of “preponderance of the evidence.” This standard requires the party bearing the burden of proof to demonstrate that their claim is more likely true than not. From a probabilistic perspective, it is often understood as requiring a demonstration of greater than 50% likelihood that the claim is true (McCauliff, 1982; Stein, 2005a), particularly in binary situations where the question is whether an assertion is true or false (Cheng, 2013). In criminal cases, the relevant standard is “proof beyond a reasonable doubt,” which is commonly understood to require a significantly higher probability (McCauliff, 1982). Probabilistic approaches to fact-finding have been extensively explored in the literature over many years (Eggleston, 1989; Cullison, 1969; Broun and Kelly, 1970; Finkelstein and Fairley, 1970; Cohen, 1977; Eggleston, 1978; McCauliff, 1982) and, despite considerable criticisms (Tribe, 1971; Allen and Stein, 2013), have remained influential, even if primarily at a conceptual level.
@@
Building upon this foundation, the essential structure of copyright disputes can be formalized by representing adjudicators’ degree of belief in terms of subjective probabilities, utilizing a discrete probability space $(\Omega,\mathcal{F},P)$ and random variables defined on this space. The condition for copyright infringement to exist can be expressed using random variables as $Copy=1\wedge Substantial=1$. Here, $Copy$ is a binary random variable representing the presence ($Copy=1$) or non-presence ($Copy=0$) of factual copying of the plaintiff’s copyrighted work during the creation of the defendant’s work. Similarly, $Substantial$ represents the presence or non-presence of substantial similarity between the two works. Typically, adjudicators can determine whether substantial similarity exists, or the truth value of $Substantial$, with certainty by directly comparing the original work to the defendant’s work presented at trial. In contrast, the absolute truth about whether the defendant engaged in factual copying remains unknown during the trial, necessitating a probabilistic approach to account for this uncertainty. The plaintiff must present sufficient evidence for the adjudicators to conclude that the probability of $Copy=1$ exceeds a certain threshold, $\lambda$, which could be 0.5, aligning with the preponderance of evidence standard.
The case law establishing that probative similarity combined with access typically suffices to prove copying can be represented in probabilistic terms as follows:
$\begin{gathered}P(Copy=1|Access=1,\,Probative=1)=\eta\\ \textrm{such that}\,\,\eta>\lambda\end{gathered}$
@@
where $Access$ and $Probative$ are binary random variables representing whether access by the defendant to the plaintiff’s original work was present and whether the original and defendant’s works are probatively similar, respectively. Suppose the adjudicators learn from the evidence presented during the trial that the defendant indeed had access to the original work ($Access=1$) and that the defendant’s work is probatively similar to the original work ($Probative=1$). This results in the adjudicators’ degree of belief that copying occurred ($Copy=1$) surpassing the threshold $\lambda$ (since $\eta>\lambda$), thereby the standard of proof being met, as desired.
This approach can accommodate the logical relationship between factual copying and access, specifically that the defendant must have had access to the original work in order to copy from it. This relationship can be represented as:
$P(Copy=1|Access=0)=0.$
As a result of imposing this restriction on $Copy$, if it is known with certainty that the defendant did not have access to the original work, the degree of belief that the defendant copied the original work must be zero, as expected.
A further step can be taken to represent the evidentiary rule that striking similarity is typically sufficient to establish copying as follows:
$\begin{gathered}P(Copy=1|Access=1,\,Striking=1)=\rho\\ P(Access=1|Striking=1)=\sigma\\ \textrm{such that}\,\,\rho\sigma>\lambda\end{gathered}$
@@
where $Striking$ is a binary random variable corresponding to the presence or non-presence of striking similarity. Assuming that the adjudicators learn during the trial that $Striking=1$ with certainty, taking into account the logical relationship between access and copying, the degree of belief about copying can be computed as:
$\begin{split}&P\left(Copy=1|Striking=1\right)\\ &=P\left(Copy=1|Striking=1,\,Access=1\right)P\left(Access=1|Striking=1\right)\\ &\,\,+P\left(Copy=1|Striking=1,\,Access=0\right)P\left(Access=0|Striking=1\right)\\ &=P\left(Copy=1|Striking=1,\,Access=1\right)P\left(Access=1|Striking=1\right)\\ &\,\,+0\cdot P\left(Access=0|Striking=1\right)\\ &=\rho\sigma>\lambda.\end{split}$
Therefore, the degree of belief of the adjudicators reaches the threshold corresponding to the applicable standard of proof, allowing them to conclude that copying has occurred.
@@ 3.3. The Inverse Ratio Rule
The so-called “inverse ratio rule,” adopted by some courts, dictates a specific relationship between proof of access and similarity: stronger proof of access lowers the level of similarity required to establish copyright infringement, while a higher level of similarity reduces the burden of proof regarding access (Aronoff, 2007; Balganesh and Menell, 2022). Although the rule has been applied for nearly a century, serious objections have been raised to its application (Aronoff, 2007). The controversial status of the inverse ratio rule gained renewed attention with its recent abrogation by the Ninth Circuit Court, [^6] Skidmore v. Led Zeppelin, 952 F.3d 1051, 1066-69 (9th Cir. 2020) (en banc). while it remains in use in some other courts (Balganesh and Menell, 2022).
The controversy surrounding the inverse ratio rule partly arises from a likely inadvertent divergence in the understanding of its scope and content. Some courts have applied the inverse ratio rule in a way that strong evidence of access lowers the level of similarity required to establish the improper appropriation prong necessary for proving infringement (Balganesh and Menell, 2022; Aronoff, 2007). However, there appears to be little justification for the inverse ratio rule operating in this manner (Balganesh and Menell, 2022). Another, more promising formulation of the inverse ratio rule pertains to the factual copying prong, where the rule adjusts the degree of similarity or the strength of evidence of access required to establish factual copying. Specifically, when strong evidence of access is present, a lower degree of similarity is sufficient to prove factual copying, whereas weak evidence of access necessitates a higher level of similarity. Conversely, a high degree of similarity lessens the evidence of access required, and vice versa. Proponents argue that the inverse ratio rule of this form follows from general principles of evidential inference (Balganesh and Menell, 2022). However, critics contend that even the inverse ration rule formulated in this way is flawed, as access is merely a necessary condition for copying, and better evidence of access does not increase the probability of copying (Aronoff, 2007). This formulation of the inverse ratio rule was also the one rejected by the Ninth Circuit Court. [^7] Id. at 1066-68.
The analysis here precisely describes the content of the inverse ratio rule as it pertains to the factual copying element in formal language and proves that the rule, defined in this way, must hold. The proof relies solely on assumptions that are naturally deduced from well-established jurisprudence. In particular, we consider a set of random variables that encapsulates the evidentiary principles discussed in the preceding section in a more generalized form (Definition 3.1).
@@
$C$ and $A$ are abbreviations of $Copy$ and $Access$, respectively, from earlier discussions, which enjoy a logical relationship (Definition 3.1 (1)). Possible values of $S$ represent varying degrees of similarity between the original and defendant’s works in the probative sense. As reflected in the well-established evidentiary rule that copying can be established by demonstrating both access and probative similarity, similarity serves as significant evidence in proving factual copying (Definition 3.1 (2)). It is reasonable, at least in principle (assuming no other specific circumstances suggest otherwise), to think that there is a monotonic relationship among the inferential values of different degrees of similarity; that is, a higher degree of similarity more strongly infers factual copying, so that $\alpha_{i+1}\geq\alpha_{i}$ holds. Similarity also has evidentiary value in establishing access to the original work (Definition 3.1 (4)), as demonstrated by the evidentiary rule regarding striking similarity. Even critics of the inverse ratio rule acknowledge the evidentiary value of similarity for establishing access, though they suggest that the strength of the inference may be weaker than assumed by others (Aronoff, 2007). It is natural to posit that such evidentiary value for access increases monotonically with similarity in principle, leading to the assumption that $\beta_{i+1,j}\geq\beta_{i,j}$ for all $i$.
Variable $EA$ is introduced to capture the interaction between levels of similarity and evidence of access (other than similarity itself). The values it can take correspond to varying degrees of the strength of evidence of access presented in court (Definition 3.1 (3)). For example, strong evidence of access might include testimony that the original work was directly handed to the defendant or the fact that the work was publicly accessible and widely known. Conversely, the fact that the original work was kept private, with little chance of the defendant encountering it, would imply a low likelihood of access. Such evidence is assumed to have direct impact only on the inference of access and is independent of other variables given $A$ (Definition 3.1 (4)). The differing strengths of evidence naturally create an ordered variation in the inferential value of access, captured by the relationship $\beta_{i,j+1}\geq\beta_{i,j}$ for all $j$.
During court proceedings, adjudicators learn both the level of similarity between the works and the strength of evidence of access presented, corresponding to a pair of values $(i,j)$. Then, the degree of belief of the adjudicators about the presence of factual copying, $C=1$, are formed given the knowledge of $S=i$ and $EA=j$.
Under these assumptions, the inverse ratio rule as it pertains to factual copying can now be precisely articulated in formal language and its correctness proved, as summarized in Proposition 3.2 (see Appendix C for proofs).
@@
The following inverse ratio rule holds for any $\lambda\in(0,1)$:
  • If $j\geq j^{\prime}$, $\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}\leq\min\{i:P(C=1|S=i,\,EA=j^{\prime})>\lambda\}$
  • If $i\geq i^{\prime}$, $\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}\leq\min\{j:P(C=1|S=i^{\prime},\,EA=j)>\lambda\}$
where the random variables are defined as in Definition 3.1.
The fact that access constitutes a necessary condition for copying does not imply that the strength of the evidence of access cannot affect the inference of copying, as critics of the inverse ratio rule argue. Stronger evidence of access can increase the likelihood of copying because the likelihood of copying is the product of two factors: (i) the likelihood of copying given the presence of access and the level of similarity, and (ii) the likelihood of access given the level of similarity and other evidence of access. Thus, stronger evidence of access leads to an increased likelihood of copying by elevating the value of the second factor in the product. Although Proposition 3.2 shows the inverse ratio rule only with weak inequalities, which can be proved with minimal assumptions, the interplay between these two factors allows the inverse ratio rule to hold in an even stronger sense, with strict inequalities, in certain specific yet highly plausible circumstances (see Appendix A for further details).
Furthermore, although the conclusion that the inverse ratio rule holds true is accepted, the analysis indicates that the reasoning posited by its proponents does not fully capture the intricacies involved. The argument previously put forward to support the inverse ratio rule contends that two pieces of relatively weak evidence—pertaining to access and similarity, respectively—possess strong inferential value when combined (Balganesh and Menell, 2022). However, the reasoning of the results concerning the inverse ratio rule demonstrates that the level of similarity and the strength of evidence on access interact also in a complementary manner, rather than merely cumulatively, due to their multiplicative effect on the probability of copying. Specifically, a higher degree of similarity increases factor (i), and even when combined with a relatively lower factor (ii) resulting from weaker evidence of access, the product of these probabilities can still surpass the necessary threshold for establishing factual copying, and vice versa. Since the product of two probabilities is always smaller than either individual factor, combining small factors—stemming from weak evidence in both similarity and access—cannot yield a stronger inference on its own. This demonstrates that the cumulative inference logic is insufficient to fully explain the interaction between evidence of access and similarity.
One might notice the product of the two probability terms is reminiscent of the so-called conjunction problem that has been a point of debate between legal scholars who support probabilistic views of legal fact-finding and those who are skeptical about these views. However, as elaborated in Appendix B, the situation at hand is different from typical scenarios that are considered in the context of the conjunction problem, and discussions surrounding the problem does not affect the analysis of the inverse ratio rule.
4. Copyright Safety for Generative AI
4.1. Evidentiary Considerations in Copyright Disputes Involving AI-Generated Works

### v5

@@ Abstract.
This paper presents a probabilistic approach to analyzing copyright infringement disputes. Evidentiary principles shaped by case law are formalized in probabilistic terms, and the “inverse ratio rule”—a controversial legal doctrine adopted by some courts—is examined. Although this rule has faced significant criticism, a formal proof demonstrates its validity, provided it is properly defined. The probabilistic approach is further employed to study the copyright safety of generative AI. Specifically, the Near Access-Free (NAF) condition, previously proposed as a strategy for mitigating the heightened copyright infringement risks of generative AI, is evaluated. The analysis reveals limitations in its justifiability and efficacy.
@@ 1. Introduction
This article employs a probabilistic approach to analyze copyright infringement disputes, seeking to bring clarity to its complexities. Based on the view that the legal fact-finding process can be understood to involve the evaluation of the probabilities that the allegations are true or false based on available evidence (Cullison, 1969; Finkelstein and Fairley, 1970; Stein, 2005), the core evidentiary principles of copyright jurisprudence are formalized in probabilistic terms. This allows for a mathematical analysis of a controversial copyright legal doctrine, the “inverse ratio rule,” whose validity has been the subject of considerable debate (Aronoff, 2007; Balganesh and Menell, 2022). While a previous study (Balganesh and Menell, 2022) notes a link between the probabilistic nature of legal fact‑finding and the inverse ratio rule, no formal probabilistic treatment of this rule appears in the literature. This paper fills that gap by offering a formal analysis of the inverse ratio rule, demonstrating that the rule must hold under natural assumptions when its content and scope are carefully delineated.
The same approach is further utilized to evaluate the Near Access-Free (NAF) condition (Vyas et al., 2023), which has been proposed to mitigate the copyright infringement risks posed by generative AI. While the authors who introduced the NAF condition left its precise legal characterization an open question (Vyas et al., 2023), subsequent studies have examined the NAF condition from a legal perspective (Lee et al., 2024; Elkin-Koren et al., 2024). This paper advances the discussion by revealing that, in certain contexts, the NAF condition operates in an ethically questionable way—reducing transparency about the materials generative AI accessed during training—and that its effectiveness remains limited.
2. Probabilistic Analysis of Copyright Disputes
@@ 2.1. Formalization of Evidentiary Principles
To prevail in a copyright infringement dispute, the copyright owner, as the plaintiff, must demonstrate two elements under the traditional Arnstein test: “(a) that defendant copied from plaintiff’s copyrighted work and (b) that the copying (assuming it to be proved) went so far as to constitute improper appropriation.” [^1] Arnstein v. Porter, 154 F.2d 464, 468 (2d Cir. 1946). These two requirements are often referred to simply as “factual copying” and “improper appropriation,” respectively (Balganesh and Menell, 2022).
Direct evidence of factual copying includes admission by the defendant or an eyewitness account of the replication or appropriative behavior, but such evidence is hard to obtain. [^2] See, e.g., Smith v. Jackson, 84 F.3d 1213, 1218 (9th Cir. 1996), stating “direct evidence of copying is not available in most cases.” Therefore, factual copying is almost invariably proved by circumstantial evidence (Gorman, 2006; Balganesh and Menell, 2022), and the evidence generally accepted by courts consists of that pertaining to access and similarity. Specifically, copying is proved based on (i) evidence showing that the defendant had access to the original work and (ii) the similarity between the original work and the defendant’s work (Balganesh and Menell, 2022; Loren and Reese, 2019). Courts generally hold that proof of access combined with a certain level of similarity, that is probative of copying, is sufficient to establish factual copying (Latman, 1990; Loren and Reese, 2019). Access pertains to the antecedent of factual copying, in the sense that the defendant having had access to the original work is a necessary condition for engaging in copying (Balganesh and Menell, 2022; Latman, 1990). In contrast, similarity emerges as a result of copying (Balganesh and Menell, 2022; Latman, 1990). When the similarity between the two works is so pronounced as to be deemed “striking,” this alone suffices to establish that copying has occurred. [^3] Arnstein, 154 F.2d at 468-69. In such cases, it can be inferred that the defendant had access to the original work based on the high degree of similarity. [^4] Ty, Inc. v. GMA Accessories, Inc., 132 F.3d 1167, 1170 (7th Cir. 1997), stating “a similarity that is so close as to be highly unlikely to have been an accident of independent creation is evidence of access.”
The second requirement, improper appropriation, is established by demonstrating that the extent of copying reaches the threshold at which the similarity between the original work and the defendant’s work can be deemed “substantial” in degree. The determination of this requirement involves a normative judgment, taking into account both quantitative and qualitative aspects of copying (Gorman, 2006; Balganesh, 2012). Thus, similarity serves a dual role: It supports the finding of factual copying and acts as an independent criterion for establishing improper appropriation (Latman, 1990; Balganesh and Menell, 2022), though the sense of similarity differs in each context, even if there may be some overlap (Loren and Reese, 2019).
@@
In court proceedings, adjudicators draw on knowledge of the frequencies of occurrences of certain events and patterns derived from experience to evaluate the likelihood of the particular circumstances in question (Stein, 2005). Although the absolute truth typically cannot be attained, the final decision in most civil cases is made based on the “preponderance of evidence.” This standard requires the party bearing the burden of proof to demonstrate that their claim is more likely true than not. From a probabilistic perspective, it is often understood as requiring a demonstration of a likelihood greater than 50% that the claim is true (McCauliff, 1982; Stein, 2005), particularly in binary situations where the question is whether an assertion is true or false (Cheng, 2013).
@@
Building on this view and considering a discrete probability space $(\Omega,\mathcal{F},P)$, essential copyright evidentiary principles can be formalized in probabilistic terms. The condition for copyright infringement to exist can be expressed using random variables as $Copy=1\wedge Substantial=1$. Here, $Copy$ is a binary random variable representing the presence ($Copy=1$) or absence ($Copy=0$) of factual copying of the plaintiff’s copyrighted work during the creation of the defendant’s work. Similarly, $Substantial$ represents the presence or absence of substantial similarity between the two works. Typically, adjudicators can determine whether substantial similarity exists, or the truth value of $Substantial$, with certainty by directly comparing the original work to the defendant’s work presented at trial. In contrast, the absolute truth about whether the defendant engaged in factual copying remains unknown during the trial, and the plaintiff must present sufficient evidence for the adjudicators to conclude that the probability of $Copy=1$ exceeds a certain threshold, $\lambda$, corresponding to the preponderance of evidence standard.
@@
The case law establishing that probative similarity combined with access typically suffices to prove copying can be represented in probabilistic terms as $P(Copy=1|Access=1,\,Probative=1)=\eta$ where $\eta>\lambda$. $Access$ and $Probative$ are binary random variables representing whether access by the defendant to the plaintiff’s original work was present and whether the original and defendant’s works are probatively similar, respectively. Suppose the adjudicators learn from the evidence presented during the trial that the defendant indeed had access to the original work ($Access=1$) and that the defendant’s work is probatively similar to the original work ($Probative=1$). This results in the adjudicators’ degree of belief that copying occurred ($Copy=1$) surpassing the threshold $\lambda$ (since $\eta>\lambda$), thus meeting the standard of proof.
@@
This approach can accommodate the aforementioned logical relationship between factual copying and access, which can be represented as $P(Copy=1|Access=0)=0$. A further step can be taken to represent the evidentiary rule that striking similarity is typically sufficient to establish copying by stipulating $P(Copy=1|Access=1,Striking=1)=\rho$ and $P(Access=1|Striking=1)=\sigma$ where $\rho\sigma>\lambda$. $Striking$ is a binary random variable corresponding to the presence or absence of striking similarity. Assuming that the adjudicators learn during the trial that $Striking=1$ with certainty, taking into account the logical relationship between access and copying, the degree of belief about copying exceeds the threshold, as desired: $P\left(Copy=1|Striking=1\right)=P\left(Copy=1|Striking=1,Access=1\right)P\left(Access=1|Striking=1\right)+0\cdot P\left(Access=0|Striking=1\right)=\rho\sigma>\lambda$.
@@ 2.2. The Inverse Ratio Rule
According to the inverse ratio rule, a stronger proof of access lowers the level of similarity required to establish copyright infringement, while a higher level of similarity reduces the burden of proof regarding access (Aronoff, 2007; Balganesh and Menell, 2022). Although the rule has been applied for nearly a century, serious objections have been raised to its application (Aronoff, 2007). The controversial status of the inverse ratio rule gained renewed attention with its recent abrogation by the Ninth Circuit Court. [^5] Skidmore v. Led Zeppelin, 952 F.3d 1051, 1066-69 (9th Cir. 2020) (en banc).
Some courts have applied the inverse ratio rule to ease the substantial similarity threshold for proving improper appropriation when access is strongly established (Balganesh and Menell, 2022; Aronoff, 2007), but such application of the rule lacks clear justification (Balganesh and Menell, 2022). A more promising formulation of the inverse ratio rule concerns the factual copying prong: Strong evidence of access lowers the similarity needed to prove factual copying, while weak evidence of access requires greater similarity; conversely, a high degree of similarity can compensate for limited evidence of access (Balganesh and Menell, 2022). In fact, this version of the inverse ratio rule can be proved by considering a set of random variables (Definition 2.1) that encapsulates the evidentiary principles discussed in the preceding section in a more generalized form.
@@
$C$ and $A$ are abbreviations of $Copy$ and $Access$, respectively, from earlier discussions, which enjoy a logical relationship (Definition 2.1 (1)). The value of $S$ represents the level of similarity in the probative sense, which serves as evidence to prove factual copying (Definition 2.1 (2)). It is reasonable, at least in principle (assuming that no other specific circumstances suggest otherwise), to think that there is a monotonic relationship among the inferential values of different degrees of similarity; that is, a higher degree of similarity more strongly infers factual copying, so $\alpha_{i+1}\geq\alpha_{i}$ holds. Similarity also has evidentiary value in establishing access to the original work (Definition 2.1 (4)), as demonstrated by the evidentiary rule regarding striking similarity. It is natural to posit that such evidentiary value for access increases monotonically with similarity in principle, leading to the assumption that $\beta_{i+1,j}\geq\beta_{i,j}$ for all $i$.
Variable $EA$ is introduced to capture the interaction between similarity and evidence of access (other than similarity itself). The values it can take correspond to varying degrees of the strength of evidence of access presented in court (Definition 2.1 (3)). For example, strong evidence of access might include testimony that the original work was directly handed to the defendant or the fact that the work was publicly accessible and widely known. Conversely, the fact that the original work was kept private, with little chance of the defendant encountering it, would imply a low likelihood of access. Such evidence is assumed to directly impact only the inference of access as specified in Definition 2.1 (4), which states that $C$ and $S$ are independent of $EA$ conditional on $A$. The differing strengths of evidence naturally create an ordered variation in the inferential value, captured by the relationship $\beta_{i,j+1}\geq\beta_{i,j}$ for all $j$.
During court proceedings, adjudicators learn both the level of similarity between the works and the strength of evidence of access presented, corresponding to a pair of values $(i,j)$. Then, the degree of belief of the adjudicators about the presence of factual copying, $C=1$, is formed given the knowledge of $S=i$ and $EA=j$.
Under these assumptions, the inverse ratio rule as it pertains to factual copying can now be precisely articulated in formal language and its correctness proved, as summarized in Proposition 2.2 (see Appendix A for a stronger version of the inverse ratio rule that holds with strict inequalities). Proofs are shown in the Appendix.
@@
The following inverse ratio rule holds for any $\lambda\in(0,1)$: If $j\geq j^{\prime}$, then $\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}\leq\min\{i:P(C=1|S=i,\,EA=j^{\prime})>\lambda\}$. Similarly, if $i\geq i^{\prime}$, then $\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}\leq\min\{j:P(C=1|S=i^{\prime},\,EA=j)>\lambda\}$.
Even though access is merely a necessary condition for copying, the strength of its evidence still affects the inference of copying by lowering the required similarity, despite critics’ argument that the inverse ratio rule improperly lowers the similarity threshold (Aronoff, 2007). This is because the likelihood of copying is the product of two factors, (i) the likelihood of copying given the presence of access and the level of similarity and (ii) the likelihood of access given the level of similarity and other evidence of access, and stronger (weaker) evidence of access enlarges (dwindles) the second factor.
Furthermore, the justification offered by the proponents of the inverse ratio rule is clarified in this paper; it is argued that moderate evidence of access and similarity can jointly support an inference of copying (Balganesh and Menell, 2022). However, since the product of two probabilities is no greater than either individual factor, combining small factors, stemming from weak evidence in both similarity and access, cannot yield a stronger inference. Instead, similarity and access work complementarily rather than cumulatively: Stronger evidence of access raises factor (ii), and even when factor (i) is modest due to a somewhat low level of similarity, their product can still clear the threshold for factual copying, and vice versa.
3. Near Access-Freeness as a Framework for Copyright Safety of Generative AI